OpenStudy (anonymous):

If an apple is hanging from a string and three flies land on it, find the probability that all three are on points that are within the same hemisphere. I am lost on how to figure this out.

OpenStudy (anonymous):

chance of first one landing on a specific hemisphere = 1/2 same foe second and third each of these probabilities are independent so the probabililies are multiplied P(3 on same hemi) = 1/2 * 1/2 * 1/2 = 1/8

OpenStudy (anonymous):

The answer is .25 but not sure how to figure this. the answer you gave is .125

OpenStudy (anonymous):

its a long time since i looked at probability theory so i could be wrong - i'll have to recheck

OpenStudy (anonymous):

ok - the answer is 0.25 because you dont include the first fly to land is not included the calculation. if the question referred to the 'left' or 'right' hemisphere then it would be 1/8 - the first fly to land can go to any hemisphere - it does not choose any specific one - but the second and third do. so its 1/2 * 1/2 = 0.25

OpenStudy (anonymous):

ok with that ?

OpenStudy (anonymous):

Awesome, thanks!!!